# Optimizing Claim Decomposition with Claim Verification Confidence

### v1

@@ Abstract
Figure 1: Left: overall framework of Decompose-Then-Verify paradigm. We define each decomposer and verifier as a LLM paired with a corresponding policy. Our dynamic decomposition is compatible with existing fact-checking systems and requires training only a decomposition policy with 4.73M parameters. Right: the figure (upper right) shows that the verification confidence of the verifier (i.e., Inst-Llama-7B with a retrieval verification policy) peaks at atomicity 1. An atomicity of -1 denotes the claim is partially trivial and tautological. The example (lower right) shows that the decomposition policy from FActScore Min et al. (2023) fails to generate subclaims that best evoke the verifier’s performance, leading to suboptimal results.
@@ 3.2 Models
Following Min et al. (2023), we determine factual labels by comparing the conditional probability of True and False from the verifier. We experiment the following three verification LLMs: a T5-3B (Raffel et al., 2023) fine-tuned on FActScore for the factuality verification task (FT-T5-3B), a Llama-7B (Touvron et al., 2023) trained on Super Natural Instructions (Wang et al., 2022) (Inst-Llama-7B), and a pretrained and instruction tuned Llama3-8B model (Meta, 2024) (Llama3-Int-8B). We employ three verification policies, each utilizing differently constructed prompts to assist verification: (1) Retrieval retrieves relevant passages from a database as evidence; (2) In-Context Example provides verification demonstrations to instruct verification; [^6] In-context examples can be found in Appendix D. and (3) No-Context directly asks the verifier for predictions, as inspired by Kadavath et al. (2022).
@@ 2 Data Source: ChatGPT
Table 1: Comparison of our DyDecomp over baselines on the test dataset. Each metric is scaled from 0 to 1. $\uparrow$ indicates higher values are preferred. We employ decomposition LLM Llama3-Inst-70B and verification LLM Llama3-Int-8B. DyDecomp consistently outperforms on atomicity 1 and 2, achieving an average improvement of 0.07 in verification confidence across two datasets and three verifiers, and a 0.12 average improvement in verification accuracy for claims sourced from PerplexityAI.
@@ 4.1 Effect of Subclaim Atomicity on Verification
Our experiment finds a strong correlation between verification confidence and accuracy across various conditions (Figure 2), including different datasets, atomiticity levels, and verifies. This finding supports verification confidence as a reliable metric for evaluating the impact of decomposition policies and as an ideal signal funneling back to policy.
@@ 5.1 Decompose-Then-Verify Paradigms
Existing factuality evaluation works have proposed various decomposition prompts revealing different characteristics of textual decomposition regarding precision (Min et al., 2023), verifiability (Song et al., 2024), coverage (Wanner et al., 2024), and atomicity (Stacey et al., 2024). More recently, research has found that superficially fine-grained subclaims with trivial information could easily inflate verification precision (Jiang et al., 2024), and decomposition benefits weaker verifiers more than stronger verifiers by generating simpler subclaims (Hu et al., 2024). Our dynamic decomposition, therefore, serves as an ideal solution to tackle the over-decomposition problem.
@@ 6 Conclusion
While dynamic decomposition addresses the problem of misalignment between decomposer and verifier, we focus on the aspect of information density (i.e., atomicity). Although well-structured, self-contained, and verifiable subclaims could further improve verification, these aspects are beyond the scope of this paper. Future research could investigate other key characteristics of decomposition and explore how to amplify their positive effects on verification through dynamic decomposition.
@@ Appendix A Limitations of Static Decomposition Policies
Our trained policy addresses the first limitation by dynamically decomposing claims and tackles the second one by leveraging verifier feedback to learn its preferred atomicities.
@@ B.1 Dataset Preparation
B.2 Training Initialization
@@ Appendix C Experiment Results
Figure 5: The verification sensitivity of dynamic decomposition as the training data size changes. The five figures (from left to right) represent claims with atomicity in the range $[0,4]$, evaluated under DyDecomp policy trained on different dataset sizes. The horizontal dashed line denotes verification confidence for original claims without decomposition. We use decomposition LLM Llama3-Inst-70B and verification LLM Llama3-Int-8B with retrieval verification policy.
@@ 2 Data Source: PerplexityAI
Table 5: Comparison of our DyDecomp over baselines on the testing set. Each metric is scaled from 0 to 1. $\uparrow$ indicates higher values are preferred. We employ decomposition LLM DeepSeek-V3 and verification LLM Llama3-Int-8B. DyDecomp consistently outperforms baselines on verification accuracy across various verifiers and input claim atomicities.

### v2

# Optimizing Decomposition for Optimal Claim Verification

@@ Abstract
Figure 1: Left: overall framework of Decompose-Then-Verify paradigm. We define each decomposer and verifier as a LLM paired with a corresponding policy. Our dynamic decomposition is compatible with existing fact-checking systems and requires training only a decomposition policy with 4.73M parameters. Right: the figure (upper right) shows that the verification confidence of the verifier (i.e., Inst-Llama-7B with a retrieval verification policy) peaks at atomicity 1. An atomicity of -1 denotes the claim is partially trivial and tautological. The example (lower right) shows that the decomposition policy from FActScore Min et al. (2023) fails to generate subclaims that best evoke the verifier’s performance, leading to suboptimal results. We provide an additional example in Appendix E to show the limitation of existing decomposition policies.
@@ 3.2 Models
Following Min et al. (2023), we determine factual labels by comparing the conditional probability of True and False from the verifier. We experiment the following three verification LLMs: a T5-3B (Raffel et al., 2023) fine-tuned on FActScore for the factuality verification task (FT-T5-3B), a Llama-7B (Touvron et al., 2023) trained on Super Natural Instructions (Wang et al., 2022) (Inst-Llama-7B), and a pretrained and instruction tuned Llama3-8B model (Meta, 2024) (Llama3-Inst-8B). We employ three verification policies, each utilizing differently constructed prompts to assist verification: (1) Retrieval retrieves relevant passages from a database as evidence; (2) In-Context Example provides verification demonstrations to instruct verification; [^6] In-context examples can be found in Appendix D. and (3) No-Context directly asks the verifier for predictions, as inspired by Kadavath et al. (2022).
@@ 2 Data Source: ChatGPT
Table 1: Comparison of our DyDecomp over baselines on the test dataset. Each metric is scaled from 0 to 1. $\uparrow$ indicates higher values are preferred. We employ decomposition LLM Llama3-Inst-70B and verification LLM Llama3-Inst-8B. DyDecomp consistently outperforms on atomicity 1 and 2, achieving an average improvement of 0.07 in verification confidence across two datasets and three verifiers, and a 0.12 average improvement in verification accuracy for claims sourced from PerplexityAI.
@@ 4.1 Effect of Subclaim Atomicity on Verification
Our experiment finds a strong correlation between verification confidence and accuracy across various conditions (Figure 2), including different datasets, atomiticity levels, and verifies. Additionally, verficaition confidence is more accessible than accuracy as it does not require ground-truth labels. These properties support verification confidence as a reliable metric for evaluating the impact of decomposition policies and as an ideal signal funneling back to policy.
@@ 4.2 Dynamic Decomposition Results
4.3 Case Study
A critical question emerges when adopting verification confidence as a proxy for accuracy: whether the policy would over-optimize for confidence in ways that did not actually improve accuracy? Our findings indicate that this phenomenon does happen, primarily due to inherent limitations in verifier capabilities. As shown in Table 1, on the PerplexityAI dataset, we observe aligned improvements in both verification confidence and actual accuracy. However, the ChatGPT dataset demonstrates substantial gains in verifier confidence with only negligible improvements in accuracy.
We posit that the disparity stems from fundamental differences in verifier capability relative to data complexity. Specifically, we find claims in the ChatGPT dataset are generally more complex and less grounded than those in the PerplexityAI dataset, making them inherently more difficult to verify. In contrast, PerplexityAI claims are typically more concise and feature explicit reference through indexed citations, enabling even less capable verifiers (Llama3-Inst-8B in this case) to evaluate factuality with greater confidence and accuracy. We provide examples in Appendix B.2 to illustrate this difference.
Our observations indicate that verifier capability is a pivotal factor in the success of dynamic decomposition, exemplifying a “bucket effect” where the weakest component constrains system performance. When paired with a sufficiently strong verifier, the dynamic decomposition can optimize the entire system in tandem, ensuring that confidence improvements translate into genuine accuracy gains.
@@ 5.1 Decompose-Then-Verify Paradigms
Existing factuality evaluation works have proposed various decomposition prompts revealing different characteristics of textual decomposition regarding precision (Min et al., 2023), verifiability (Song et al., 2024), coverage (Wanner et al., 2024), and atomicity (Stacey et al., 2024). More recently, research has found that superficially fine-grained subclaims with trivial information could easily inflate verification precision (Jiang et al., 2024), and decomposition benefits weaker verifiers more than stronger verifiers by generating simpler subclaims (Hu et al., 2024). Our dynamic decomposition effectively addresses the over-optimization problem by providing necessary controllability over the decomposition process.
@@ 6 Conclusion
While dynamic decomposition addresses the problem of misalignment between decomposer and verifier, we focus on the aspect of information density (i.e., atomicity). Although well-structured, self-contained, and verifiable subclaims could further improve verification, these aspects are beyond the scope of this paper. Future research could investigate other key characteristics of decomposition and explore how to amplify their positive effects on verification through dynamic decomposition (e.g., using a more powerful decomposer).
@@ Appendix A Limitations of Static Decomposition Policies
Our dynamic decomposition is the first solution to tackle these limitations by explicitly exploring various decomposition complexities during training. As a result, it achieves a more generalizable performance and consistently outperforms static methods across varying input atomicities.
@@ B.1 Data Preparation
B.2 Data Example
Data from ChatGPT Quintus Sosius Senecio was a Roman senator and military commander who lived during the 2nd century AD. He was a member of the senatorial class and served in various high-ranking positions in the Roman government, including as consul in 128 AD.
In addition to his political career, Quintus Sosius Senecio was also a distinguished military leader. He served as a legate (commander) in several provinces, including Africa and Pannonia, and was awarded triumphal honors for his victories over the enemies of Rome.
Quintus Sosius Senecio was also known for his literary works, including a history of the Roman Empire, which unfortunately has not survived to the present day.
Overall, Quintus Sosius Senecio was a prominent figure in Roman society, known for his military prowess, political acumen, and intellectual pursuits.
Data from PerplexityAI Quintus Sosius Senecio was a Roman consul who lived during the 1st and 2nd centuries AD[1]. He was married to Julia, the daughter of Frontinus, a prominent Roman civil engineer, author, soldier, and senator[3]. Quintus Sosius Senecio was the father of Sosia Polla[1], who married Quintus Pompeius Falco, a consul in AD 109[4].
Quintus Sosius Senecio was a friend of Plutarch, a Greek Middle Platonist philosopher, historian, biographer, essayist[2]. Plutarch wrote about several Roman nobles in his works including Quintus Sosius Senecio and Titus Avidius Quietus[2].
B.3 Training Initialization
@@ Appendix C Experiment Results
Figure 5: The verification sensitivity of dynamic decomposition as the training data size changes. The five figures (from left to right) represent claims with atomicity in the range $[0,4]$, evaluated under DyDecomp policy trained on different dataset sizes. The horizontal dashed line denotes verification confidence for original claims without decomposition. We use decomposition LLM Llama3-Inst-70B and verification LLM Llama3-Inst-8B with retrieval verification policy.
@@ 2 Data Source: PerplexityAI
Table 5: Comparison of our DyDecomp over baselines on the testing set. Each metric is scaled from 0 to 1. $\uparrow$ indicates higher values are preferred. We employ decomposition LLM DeepSeek-V3 and verification LLM Llama3-Inst-8B. DyDecomp consistently outperforms baselines on verification accuracy across various verifiers and input claim atomicities.
@@ Appendix D Prompts
Appendix E Example
Method Results
Original After retiring from tennis in 1996, Sabatini pursued other interests, including fashion and fragrance design. She also established the Gabriela Sabatini Foundation, which supports children’s education and sports programs in Argentina.
[…]
Table 6: Example of decomposition results from different methods, where we use DeepSeek-V3 as the decomposer. Our DyDecomp policy is trained for the verifier Llama3-Inst-8B with No-Context policy.
We provide an example in Table 6 to show the differences between subclaims decomposed using our DyDecomp policy and those derived from popular decomposition methods. For the verifier Llama3-Inst-8B using No-Context verification policy, which according to Figure 4 has the best verification performance at atomicity level 2 (i.e., input contains 4 pieces of atomic information). Evidently, only our trained DyDecomp policy successfully generates subclaims that closely match the verifier’s preferred atomicity level. These examples highlight how existing decomposition methods can lead to suboptimal performance in fact-checking systems.